# SpineBench: A Clinically Salient, Level-Aware Benchmark Powered by the SpineMed-450k Corpus

### v2

@@ Appendix B Contributions and Acknowledgments
    Ming Zhao
    $\pi^{3}$ Lab, Jilin University
@@
    Wenhui Dong
    $\pi^{3}$ Lab, Nanjing University
@@
    Yang Zhang
    The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Xiang Zheng
    $\pi^{3}$ Lab, Institute of Automation, Chinese Academy of Sciences
@@
    Zhonghao Zhang
    $\pi^{3}$ Lab, Ningxia University
@@
    Zi An Zhou
    $\pi^{3}$ Lab, Zhejiang University
@@
    Yunzhi Guan
    Huashan Hospital, Fudan University
@@
    Liukun Xu
    The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Zhaoyang Gong
    Huashan Hospital, Fudan University
@@
    Zhicheng Zhang
    The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Dachuan Li
    Huashan Hospital, Fudan University
@@
    Xiaosheng Ma
    Huashan Hospital, Fudan University
@@
    Yuli Ma
    Sanyou Medical
@@
    Changjiang Jiang
    Wuhan University
@@
    Lixia Tian
    Beijing Jiaotong University
@@
    Qixin Chen
    The Second Affiliated Hospital, Zhejiang University
@@
    Kaishun Xia
    The Second Affiliated Hospital, Zhejiang University
@@
    Pingping Liu
    Jilin University
@@
    Tongshun Zhang
    Jilin University
@@
    Zhiqiang Liu
    $\pi^{3}$ Lab,Huazhong University of Science and Technology
@@
    Zhongan Bi
    $\pi^{3}$ Lab,Zhejiang University
@@
    Chenyang Si
    Nanjing University
@@
    Tiansheng Sun
    The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Caifeng Shan
    Nanjing University
@@ Appendix C RELATED WORK
From Generalist Models to Domain Adaptation. Recent advances in Large Vision-Language Models (LVLMs), such as GPT-4V (OpenAI, 2023) and Gemini2.5-pro (Google, 2025a), have demonstrated significant progress in multimodal tasks (Yang et al., 2023; Team et al., 2023). However, when applied to the medical domain, their generalist nature becomes a distinct limitation. Multiple evaluations consistently show that while promising, these models lack the domain-specific expertise required for complex diagnostic tasks, performing below the level of human specialists (AlSaad et al., 2024). This inherent limitation of generalist models has fueled a clear and necessary trend toward specialization. In response, specialized medical LVLMs like LLaVA-Med (Li et al., 2023) and PMC-LLaMA (Wu et al., 2024) have been developed, fine-tuned on large biomedical corpora. Nevertheless, this approach still has shortcomings. For instance, in spinal diagnostics, a critical task is the synthesis of data from multimodal imaging—such as X-ray, CT, and MRI—to formulate a single, "level-aware" diagnosis. This integrative reasoning process, which requires localizing findings to specific vertebral levels, is a clinical skill that cannot be acquired from static, descriptive datasets alone. This further underscores a core principle: for high-stakes clinical applications, deep, narrow expertise is far more valuable than broad, superficial general knowledge. A powerful example validating this principle is OralGPT (Hao et al., 2025), a model trained on a small, highly curated dataset of intraoral photographs, which achieves performance comparable to state-of-the-art generalist models within its niche. This paradigm shift from generalist to specialist models is now clearly evident across numerous medical fields, from oncology to pathology (Qiu et al., 2023; Sarabadani et al., 2025; Yang et al., 2025; 2024; Barrit et al., 2024; Mo et al., 2025; Deng et al., 2024; Xue et al., 2024; Bhaumik et al., 2023; Na, 2024; Guo et al., 2025).
@@
Foundational Datasets and the Cognitive Gap. Progress in AI is fundamentally tied to the quality of training data. Foundational datasets like MIMIC-CXR (Johnson et al., 2019) and CheXpert (Irvin et al., 2019) have been instrumental for tasks like chest radiograph classification. Moving up in complexity are datasets for interactive Visual Question Answering (VQA). For instance, VQA-RAD (Lau et al., 2018) was manually constructed by clinicians asking naturally occurring questions about radiology images, representing a step toward more dynamic reasoning. More recently, large-scale efforts like MedTrinity-25M (Xie et al., 2024b) have emerged, providing over 25 million images with multi-granular annotations to support a wide range of tasks. Within the spine domain itself, public datasets have primarily supported foundational computer vision tasks. The VerSe dataset (Sekuboyina et al., 2021), for example, is a critical resource providing CT scans with precise voxel-level annotations for vertebral segmentation and identification. Other datasets (Lee et al., 2024a; Ibrahim et al., 2025a) have followed a similar focus, providing valuable benchmarks for segmentation of the lumbar spine from MRI. However, these resources are designed to support lower-level cognitive tasks like perception ("Where is the L4 vertebra?") or classification ("Is a fracture present?"). They do not provide the necessary data to train models for the highest level of clinical cognition: synthesizing multimodal information into a comprehensive diagnosis and treatment plan. This reveals a crucial gap between existing data and the needs of clinical practice, a gap our work aims to fill.
@@
Appendix D Training Strategy
Table 6: Training configurations across different stages.
Configuration Stage-1 Stage-2 Stage-3
Dataset PubMedVision-150k Spine-open Spine-chat
orthopedics-230k Spine-choice Spine-chat-reasoning
MedThoughts-8K Spine-report
Medical-R1-Distill-Data Spine-report-reasoning
medical-o1-reasoning-SFT
Learning Rate 1e-5 1e-5 1e-6
Max Length 16,284 16,284 49,152
DeepSpeed zero2 zero2 zero3
Epochs 1 1 1
Appendix E Performance comparison on medical report generation subtasks
@@
Appendix F Picture Context Matching Algorithm

### v3

@@ Appendix B Contributions and Acknowledgments
    Ming Zhao, $\pi^{3}$ Lab, Jilin University
@@
    Wenhui Dong, $\pi^{3}$ Lab, Nanjing University
@@
    Yang Zhang, The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Xiang Zheng, $\pi^{3}$ Lab, Institute of Automation, Chinese Academy of Sciences
@@
    Zhonghao Zhang, $\pi^{3}$ Lab, Ningxia University
@@
    Zi An Zhou, $\pi^{3}$ Lab, Zhejiang University
@@
    Yunzhi Guan, Huashan Hospital, Fudan University
@@
    Liukun Xu, The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Zhaoyang Gong, Huashan Hospital, Fudan University
@@
    Zhicheng Zhang, The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Dachuan Li, Huashan Hospital, Fudan University
@@
    Xiaosheng Ma, Huashan Hospital, Fudan University
@@
    Yuli Ma, Sanyou Medical
@@
    Changjiang Jiang, Wuhan University
@@
    Lixia Tian, Beijing Jiaotong University
@@
    Qixin Chen, The Second Affiliated Hospital, Zhejiang University
@@
    Kaishun Xia, The Second Affiliated Hospital, Zhejiang University
@@
    Pingping Liu, Jilin University
@@
    Tongshun Zhang, Jilin University
@@
    Zhiqiang Liu, $\pi^{3}$ Lab, Huazhong University of Science and Technology
@@
    Zhongan Bi, $\pi^{3}$ Lab, Zhejiang University
@@
    Chenyang Si, Nanjing University
@@
    Tiansheng Sun, The Fourth Medical Center of People’s Liberation Army General Hospital
@@
    Caifeng Shan, Nanjing University
@@ Appendix C RELATED WORK
From Generalist Models to Domain Adaptation. Recent advances in Large Vision-Language Models (LVLMs), such as GPT-4V (OpenAI, 2023) and Gemini2.5-pro (Google, 2025b), have demonstrated significant progress in multimodal tasks including image and video(Yang et al., 2023; Team et al., 2023; Zheng et al., 2023; Jiang et al., 2025; Guo et al., 2025a; Zhu et al., 2024; Jiang et al., 2026). However, when applied to the medical domain, their generalist nature becomes a distinct limitation. Multiple evaluations consistently show that while promising, these models lack the domain-specific expertise required for complex diagnostic tasks, performing below the level of human specialists (AlSaad et al., 2024). This inherent limitation of generalist models has fueled a clear and necessary trend toward specialization. In response, specialized medical LVLMs like LLaVA-Med (Li et al., 2023) and PMC-LLaMA (Wu et al., 2024) have been developed, fine-tuned on large biomedical corpora. Nevertheless, this approach still has shortcomings. For instance, in spinal diagnostics, a critical task is the synthesis of data from multimodal imaging—such as X-ray, CT, and MRI—to formulate a single, "level-aware" diagnosis. This integrative reasoning process, which requires localizing findings to specific vertebral levels, is a clinical skill that cannot be acquired from static, descriptive datasets alone. This further underscores a core principle: for high-stakes clinical applications, deep, narrow expertise is far more valuable than broad, superficial general knowledge. A powerful example validating this principle is OralGPT (Hao et al., 2025), a model trained on a small, highly curated dataset of intraoral photographs, which achieves performance comparable to state-of-the-art generalist models within its niche. This paradigm shift from generalist to specialist models is now clearly evident across numerous medical fields, from oncology to pathology (Qiu et al., 2023; Sarabadani et al., 2025; Yang et al., 2025; 2024; Barrit et al., 2024; Mo et al., 2025; Deng et al., 2024; Xue et al., 2024; Bhaumik et al., 2023; Na, 2024; Guo et al., 2025b).
@@
Foundational Datasets and the Cognitive Gap. Progress in AI is fundamentally tied to the quality of training data. Foundational datasets like MIMIC-CXR (Johnson et al., 2019) and CheXpert (Irvin et al., 2019) have been instrumental for tasks like chest radiograph classification. Moving up in complexity are datasets for interactive Visual Question Answering (VQA). For instance, VQA-RAD (Lau et al., 2018) was manually constructed by clinicians asking naturally occurring questions about radiology images, representing a step toward more dynamic reasoning. More recently, large-scale efforts like MedTrinity-25M (Xie et al., 2024b) have emerged, providing over 25 million images with multi-granular annotations to support a wide range of tasks.
Within the spine domain itself, public datasets have primarily supported foundational computer vision tasks. As summarized in Table 1, existing datasets largely focus on single-modality perception tasks. For example, the RSNA LumbarDISC dataset (Richards et al., 2025) provides a large-scale MRI benchmark but is limited to the classification of stenosis severity grades. Similarly, BUU-LSPINE (Klinwichit et al., 2023) focuses on spondylolisthesis detection in X-rays, while the VerSe 2020 (Liebl et al., 2021) and Lumbar Spine MRI (van der Graaf et al., 2024) datasets provide voxel-level masks for segmentation tasks in CT and MRI, respectively.
However, these resources are primarily designed to support lower-level cognitive tasks like perception ("Where is the L4 vertebra?") or classification ("Is a fracture present?"). They lack the multimodal integration and instruction-following structure required to train models for the highest level of clinical cognition: synthesizing multimodal information into a comprehensive diagnosis and treatment plan. This reveals a crucial gap between existing data paradigms and the needs of clinical practice—a gap our work aims to fill by introducing the first large-scale instruction-tuning corpus designed for full-spectrum clinical reasoning.
@@
Appendix D Performance comparison on medical report generation subtasks
@@
Appendix E Picture Context Matching Algorithm
@@
Appendix F Detailed Distribution Analysis of SpineBench
To further validate the clinical representativeness and unbiased nature of SpineBench, we provide a comprehensive statistical breakdown of the testing QA set (487 items with specific subspecialty tags). These statistics demonstrate that SpineBench achieves rigorous coverage across diverse pathologies and data sources.
F.1 Disease Subspecialty Distribution
As shown in Table 10, the benchmark covers 14 distinct spinal subspecialties. Notably, high-stakes and complex conditions such as Tumor and Osteomyelitis (16.72%) and Complications (10.80%) are heavily represented, ensuring that the evaluation reflects performance on critical clinical challenges rather than being dominated by common degenerative cases.
Table 10: Detailed Distribution of Spine Subspecialties in SpineBench
Rank Subspecialty Count Percentage
1 Tumor and Osteomyelitis 96 16.72%
2 Complications 62 10.80%
3 Dysplastic and Congenital Deformities 49 8.54%
4 Idiopathic Scoliosis 47 8.19%
5 The Thoracic and Lumbar Degenerative Spine 34 5.92%
6 General Considerations 33 5.75%
7 Kyphosis and Postlaminectomy Deformities 31 5.40%
8 Anatomic Approaches 31 5.40%
9 Adult Spinal Deformity 26 4.53%
10 Spondylolisthesis 23 4.01%
11 Trauma 23 4.01%
12 Neuromuscular Spine Deformity 12 2.09%
13 Biomechanics 11 1.92%
14 The Cervical Degenerative Spine 9 1.57%
Total All Subspecialties 487 100.00%
F.2 Data Source Diversity
Table 11 illustrates the provenance of the testing data. The dataset maintains a rigorous balance between Academic Resources (45.8%), which ensure theoretical precision, and Real-world Clinical Data (54.2%), which test practical diagnostic robustness. Crucially, the Hospital Databases (33.5%) consist of private, internal cases that ensure a zero-leakage evaluation environment.
Table 11: Detailed Source Distribution of SpineBench QAs
Category Specific Composition Count % Evaluation Goal
Textbooks Standard Medical Textbooks 203 41.7% Theoretical Foundation
Hospital Databases Proprietary data related to privacy 163 33.5% Surgical Precision & Zero-Leakage
Clinical Case Reports Real-world Case Reports 101 20.7% Complex Clinical Scenarios
Literature Academic Papers 17 3.5% Cutting-edge Knowledge
Question Bank Medical Exam Questions 3 0.6% Standardized Logic
Total 487 100%